# Artificial Precision Polarization Array: Sensitivity for the axion-like dark matter with clock satellites

### v1

@@ III.1 Bayesian analysis
$\mathrm{Tr}\left({\textstyle\sum_{ii}^{(s)}}{\textstyle\sum_{ii}^{(s)}}\right)$ $=4\frac{g_{a\gamma}^{4}}{m_{a}^{4}}N_{i}^{2}\rho_{\text{DM}}^{2}\left[1-\mathrm{cos}\big(m_{a}L_{i}\big)\frac{\mathrm{sin}\big(y_{ei}\big)}{y_{ei}}\right]^{2},$
@@
$\mathrm{Tr}\left({\textstyle\sum_{ij}^{(s)}}{\textstyle\sum_{ij}^{(s)}}\right)$ $=\frac{g_{a\gamma}^{4}}{m_{a}^{4}}N_{i}N_{j}\rho_{\text{DM}}^{2}\left[1+\frac{\mathrm{sin}^{2}\big(y_{ij}\big)}{y_{ij}^{2}}\right$.
$\quad\left.+2\mathrm{cos}\big(m_{a}\big(L_{i}-L_{j}\big)\big)\frac{\mathrm{sin}\left(y_{ij}\right)}{y_{ij}}\right].$
Here $L_{i}$ is the distance between the $i$-th transmitting satellite and the receiving satellite, $y_{ij}\equiv\left|\mathbf{r}_{i}-\mathbf{r}_{j}\right|/\lambda_{de}$ is the dimensionless distance parameter and $N_{i}$ denotes the total number of pulse signals emitted by the $i$-th transmitting satellite within the coherence time $\tau_{c}$.
@@
In the numerical simulation, we assume a total of $N_{\mathrm{sat}}=6$ transmitting satellites. The receiving satellite is placed at the center, while the transmitting satellites are randomly distributed in all directions with their distances from the receiving satellite given by $3\times 10^{5}\mathcal{U}\big[0.5,1.5\big]$ $\mathrm{km}$. White noise is generated according to Eq. (III.1) and Eq. (III.2), with $\sigma_{\text{err}}=0.3^{\circ}$. For each transmitting satellite ($i=1,2,\dots,6$), the number of pulse signals is set to $N_{i}=10^{5}$. Using Eq. (III.4), we show the result $g_{95\%}$ of the numerical simulation in Fig. 2. Under the assumption that the satellite network occupies a single coherent axion domain, the parameter $y<1$ and in practice $y\ll 1$, thus $\mathrm{sin}\big(y\big)/y\simeq 1$. Since the cosine factor preceding $\mathrm{sin}\big(y\big)/y$ is bounded, Eq. (III.4) is insensitive to the pairwise distances between satellites.
@@
Figure 2: Bayesian 95% C.L. upper limit $g_{95\%}$ on the function $g_{a\gamma}\big(m_{a}\big)$, which is calculated from $\big\langle q\big\rangle=2.71$ and Eq. (III.4).
@@ III.2 Frequentist analysis-GLSP
The GLSP is a spectral analysis technique designed to detect periodic signals in irregularly sampled time series. It employs a least squares fitting procedure to scan the trial frequencies and compute the power spectrum at each frequency. Compared to traditional Fourier transform methods, the GLSP is particularly advantageous for handling unevenly sampled data, making it well suited for astronomical observations [10.3847/1538-4365/aab766]. According to [10.3847/1538-4365/aab766], the GLSP is expressed as
$P_{LS}\big(f\big)$ $=\frac{1}{\sigma^{2}_{y}\sum_{i}\omega_{i}}\Bigg\{\frac{\big[\sum_{i}\omega_{i}\big(y_{i}-\bar{y}\big)\cos\big(2\pi f\big(t_{i}-\tau\big)\big)\big]^{2}}{\sum_{i}\omega_{i}\cos^{2}\big[2\pi f\big(t_{i}-\tau\big)\big]}$ (III.5)
@@
$\tau=\frac{1}{4\pi f}\arctan\frac{\sum_{i}\omega_{i}\sin\big(4\pi ft_{i}\big)}{\sum_{i}\omega_{i}\cos\big(4\pi ft_{i}\big)},$ (III.6)
@@
$\omega_{i}$ $=\frac{1}{\sigma^{2}_{i}},\quad\bar{y}=\frac{\sum_{i}\omega_{i}y_{i}}{\sum_{i}\omega_{i}},$ (III.7)
@@
A distinctive feature of the window function in ground-based radio telescope observations is that, in order to detect faint pulsar signals, observations are restricted to nighttime. Consequently, the resulting time series is highly non-uniformity sampled. Moreover, due to the annual modulation introduced by the Earth’s orbital motion, the data are not only influenced by the ionospheric interference discussed earlier but also subject to additional deterministic periodic effects [10.3847/1538-4365/aab766]. Together, these factors give rise to pronounced sidelobes and spectral aliasing in the Lomb–Scargle periodogram [10.3847/1538-4365/aab766], thereby increasing the False Alarm Probability (FAP) and potentially obscuring the true signal peaks. As illustrated in Fig. 3, the irregular temporal sampling causes the target axion signal to be masked by numerous spurious peaks in the power spectrum. This effect becomes increasingly severe as the observational S/N decreases, and in extreme cases, the signal peak may fall entirely below the significance threshold, rendering it undetectable. The figure also demonstrates that many interference peaks approach the 10% FAP level, substantially complicating the reliable identification of genuine signals.
Figure 3: Numerical simulation: The GLSP (blue curve) derived from simulated ground-based radio telescope data. The simulated dataset contains only Gaussian white noise with $\text{S/N}=1$. The vertical dashed line marks the target signal frequency corresponding to $m_{a}=10^{-22}\mathrm{eV}$. The red horizontal line indicates the 10% (FAP) threshold. Since the peak at the target frequency lies below this threshold, the likelihood of the detected peak representing a genuine signal is less than 90%.
@@
In contrast, within our proposed framework, once the satellite network is deployed, it can operate in continuous monitoring mode. In the idealized case, the corresponding window function of space-based observations reduces to $W_{{t_{i}}}\big(t\big)=1$, thereby eliminating the non-uniformity inherent in ground-based sampling. To illustrate this, we consider a simplified scenario in which the observed signal contains only a single axion induced frequency component, $m_{a}=10^{-22}\mathrm{eV}$. The transmitting satellites emit periodic pulse signals towards the receiving satellite at a fixed interval $\Delta t=\delta t+\sigma=\big(1+\xi_{\delta t}\big)\delta t$. In addition, the system is affected by Gaussian white noise, $\phi_{\text{noise}}$. Under these assumptions, the APPA measured at the receiving end can be expressed as
$\Delta\phi_{sim}=\phi_{a}\mathrm{cos}\big(2\pi\nu+\varphi\big)+\phi_{noise}.$ (III.8)
@@
Similar to Fig. 3, we consider the case of $\text{S/N}=1$ and a target signal frequency of the axion with mass $m_{a}=10^{-22}\mathrm{eV}$. However, in this scenario, the observation epochs are distributed with high regularity over a continuous 10-year period. The corresponding simulation results are presented in Fig. 4.
@@
Figure 4: Numerical simulation: The GLSP (blue curve) derived from APPA observational data. The simulated signal includes only Gaussian white noise with S/N=1. The vertical dashed line marks the target signal frequency of the axion with mass $m_{a}=10^{-22}\mathrm{eV}$. The horizontal lines indicate FAP levels of 10% (red), 5% (yellow), and 1% (green).
@@
If the observational data are of low quality and the target frequency peak is completely masked by interference peaks, the GLSP method becomes ineffective. In such cases, for further information, the reader is referred to [10.1103/PhysRevD.111.062005, 10.1088/1475-7516/2022/06/014]. Once the GLSP method successfully extracts the target axion signal frequency $m_{a}$ from the pulsed signals received by the satellite, the APPA sensitivity curve for axion signal detection can be derived using Eq. (B.1) and Eq. (III.8).
We assume that the satellite network possesses a detection sensitivity of $0.01^{\circ}$ for linear polarization angle rotation, and that the instrumental noise is parameterized according to Eq. (III.1) and Eq. (III.2). Under these assumptions, the sensitivity curve can be obtained subject to the constraint of a total observation time of $T_{\text{obs}}\sim\mathcal{O}\big(10^{1}\big)$yr (see Fig. 5). Note that for Eq. (B.1), when condition $\mathrm{cos}\big(m_{a}T\big)\simeq 1$ is no longer satisfied, the expectation value of $\big[\big(\alpha_{o}^{2}+\alpha_{s}^{2}\big)-2\alpha_{o}\alpha_{s}\mathrm{cos}\big(m_{a}T\big)\big]^{\frac{1}{2}}$ must be evaluated individually for each value of $m_{a}$. As in the analysis of Sec. III.1, because the signal transmission time $T\equiv L/c$ enters only via a cosine factor $\mathrm{cos}\left(m_{a}T\right)$, $\phi_{a}$ also remains insensitive to the pairwise distances between satellites.
Figure 5: Detection range and sensitivity curve of the APPA satellite network. The detection sensitivity is set to $\phi_{a}=0.01^{\circ}$ in Eq. (III.8), consistent with current observational limits on linear polarization angle rotation, with minor fluctuations attributed to instrumental noise. The black dashed line marks the critical boundary beyond which the condition $\mathrm{cos}\big(m_{a}T\big)\simeq 1$ is no longer valid.
Finally, we present a comparison between the simulation results of this work and existing observational constraints in a single figure (see Fig. 6). Owing to the long observational baseline $T_{\text{obs}}$ and the highly stable signal emission cadence $\Delta t$, APPA is theoretically capable of probing an exceptionally broad range of axion induced oscillation frequencies. However, as shown in Fig. 6, although APPA theoretically possesses an extremely wide detection range, the domain where it truly demonstrates an advantage in axion signal detection is limited to $m_{a}\sim\mathcal{O}\big(10^{-22}–10^{-19}\big)$eV, both the 95% confidence upper limit $g_{95\%}$ derived from Bayesian analysis and the detection sensitivity $g_{a\gamma}$ obtained via frequentist analysis highlight the significant advantages of the APPA framework over current ground-based observations.
Figure 6: Comparison between the APPA simulation results and existing observational constraints. Color coding: red, CAST [10.1038/nphys4109]; green, PPA [2412.02229]; blue, PPTA [10.1088/1475-7516/2022/06/014]; light blue lines, simulation results from this work, the upper line represents the results of the frequentist analysis, whereas the lower line represents the results of the Bayesian analysis.
@@ IV Conclusion
Inspired by the artificial precision timing array in [2401.13668], we investigate the APPA concept, which leverages an artificial satellite network to achieve precise measurements of ultralight scalar axion signals. Compared with PPA data obtained from ground-based observatories, APPA offers several significant advantages: (i) The number of uncertain or uncontrollable physical parameters is greatly reduced, since most parameters can be explicitly determined and flexibly adjusted; (ii) Complex or periodic physical disturbances are largely eliminated, thereby reducing the burden of data analysis and enhancing the quality of the raw observational data; (iii) Thanks to the ultra precise clocks of satellites and the tunability of pulsed signal parameters, the S/N, periodicity, and long term stability of the measurements are well ensured. Consequently, APPA provides an exceptionally broad detectable frequency range.
To clarify the feasibility and advantages of APPA, we employ two complementary approaches: Bayesian analysis and frequentist analysis. Both methods exhibit insensitivity to the spatial scale of the satellite network. The network’s spatial scale can only constrain the upper bound of axion mass that APPA can probe. However, Fig. 6 shows that APPA exhibits a detection advantage only within $m_{a}\sim\mathcal{O}\big(10^{-22}–10^{-19}\big)$eV, which lies well below the detection upper bound. Therefore, the spatial distribution scale of the satellite network is not a primary concern in the detailed analysis.
For the Bayesian approach, we avoid prohibitively large scale computations by using the convenient expression in Eq. (III.4) from [10.1140/epjc/s10052-011-1554-0, 10.1103/PhysRevD.103.076018], which allows us to efficiently derive the $g_{95\%}$ constraint. Compared with traditional ground-based observations, APPA achieves a tighter $g_{95\%}$ bound, highlighting its strong detection potential. For the frequentist approach, we adopt the GLSP, a standard method for detecting periodic sinusoidal signals. MC simulations demonstrate that, even under low S/N conditions, APPA produces GLSP spectra with more pronounced target signal peaks, significantly reduced FAP, and fewer interference peaks compared with conventional observations.
@@
V Acknowledgments
This work is supported by the National Key Research and Development Program of China (Nos. 2023YFC2206200, No.2021YFC2201901), the National Natural Science Foundation of China (No.12375059), and the Project of National Astronomical Observatories, Chinese Academy of Sciences (No.E4TG6601). We thank helpful conversion with Prof. Szabolcs Marka.

### v2

@@ III.1 Likelihood analysis
$\mathrm{Tr}\left({\textstyle\sum_{ii}^{(s)}}{\textstyle\sum_{ii}^{(s)}}\right)$ $=4\frac{g_{a\gamma}^{4}}{m_{a}^{4}}N_{i}^{2}\rho_{\text{DM}}^{2}\Big[1-\mathrm{cos}\big(m_{a}L_{i}\big)\mathrm{sinc}\big(y_{ei}\big)\Big]^{2},$
@@
$\mathrm{Tr}\left({\textstyle\sum_{ij}^{(s)}}{\textstyle\sum_{ij}^{(s)}}\right)$ $=\frac{g_{a\gamma}^{4}}{m_{a}^{4}}N_{i}N_{j}\rho_{\text{DM}}^{2}\Big[1+\mathrm{sinc}^{2}\big(y_{ij}\big)$
$\quad+2\mathrm{cos}\big(m_{a}\big(L_{i}-L_{j}\big)\big)\mathrm{sinc}\big(y_{ij}\big)$
$\quad+f\big(y_{ei},y_{ej}\big)\Big].$
The factor $f\big(y_{ei},y_{ej}\big)$ in the above expression is given by
$f\big(y_{ei},y_{ej}\big)$ $=-2\Big[\mathrm{cos}\big(m_{a}L_{i}\big)\big(\mathrm{sinc}\big(y_{ei}\big)+\mathrm{sinc}\big(y_{ej}\big)\mathrm{sinc}\big(y_{ij}\big)\big)$ (III.5)
$\quad+\mathrm{cos}\big(m_{a}L_{j}\big)\big(\mathrm{sinc}\big(y_{ej}\big)+\mathrm{sinc}\big(y_{ei}\big)\mathrm{sinc}\big(y_{ij}\big)\big)\Big]$
$\quad+\Big[\mathrm{sinc}^{2}\big(y_{ei}\big)+\mathrm{sinc}^{2}\big(y_{ej}\big)$
$\quad+2\mathrm{cos}\big(m_{a}\big(L_{i}+L_{j}\big)\big)\mathrm{sinc}\big(y_{ei}\big)\mathrm{sinc}\big(y_{ej}\big)\Big].$
Here $L_{i}$ is the distance between the $i$-th transmitting satellite and the receiving satellite, $y_{ij}\equiv\left|\mathbf{r}_{i}-\mathbf{r}_{j}\right|/\lambda_{\text{de}}$ and $y_{ei}\equiv\left|\mathbf{r}_{i}\right|/\lambda_{\text{de}}$ are the dimensionless distance parameters and $N_{i}$ denotes the total number of pulse signals emitted by the $i$-th transmitting satellite within the coherence time $\tau_{c}$. $\mathrm{sinc}\big(y\big)\equiv\mathrm{sin}\big(y\big)/y$ is the sinc function. [33] notes that the factor $f\big(y_{ei},y_{ej}\big)$ is negligible when $y_{ei}\gg 1$, but must be retained when $y_{ei}\leq 1$. For pulsars located at cosmological distances, the condition $y_{ei}\gg 1$ is typically satisfied; in contrast, for the APPA considered in this work, the satellite network occupies a single coherent axion domain, the parameter $y_{ei}\leq 1$ and in practice $y_{ei}\ll 1$, thus $\mathrm{sinc}\big(y_{ei}\big)\simeq 1$. Consequently, $f\big(y_{ei},y_{ej}\big)$ plays an important role in the analysis and cannot be neglected.
@@
In the numerical simulation, we assume a total of $N_{\mathrm{sat}}=6$ transmitting satellites. The receiving satellite is placed at the center, while the transmitting satellites are randomly distributed in all directions with their distances $L_{i}$. White noise is generated according to Eq. (III.1) and Eq. (III.2), with $\sigma_{\text{err}}=0.3^{\circ}$. For each transmitting satellite ($i=1,2,\dots,6$), the number of pulse signals is set to $N_{i}=10^{5}$. Here, using Eq. (III.4) we show the impact of the axion mass $m_{a}$ and the characteristic scale $L$ (the average distance from the transmitting satellite to the receiving satellite) of the satellite network on the 95% C.L. upper limits $g_{95\%}$, as illustrated in Fig. 3 and Fig. 4.
@@
Figure 3: 95% C.L. upper limit $g_{95\%}$ on the function $g_{a\gamma}\big(m_{a}\big)$, which is calculated from $\big\langle q\big\rangle=2.71$ and Eq. (III.4). Color coding: Red, $L=2.5\times 10^{6}\mathrm{km}$ (LISA [3]); Magenta, $L=2.6\times 10^{8}\mathrm{km}$ (ASTROD-GW [37]); Pink, $L=7.5\times 10^{8}\mathrm{km}$ (Jupiter’s orbit, 5AU).
Figure 4: 95% C.L. upper limit $g_{95\%}$ on the function $g_{95\%}\big(L\big)$, which is calculated from $\big\langle q\big\rangle=2.71$ and Eq. (III.4). The transition point at which $g_{95\%}$ changes from being sensitive to insensitive to the distance is approximately given by the corresponding Compton wavelength $\lambda_{\text{com}}$ of axion. Color coding: Cyan, $m_{a}=10^{-14}\mathrm{eV}$; Green, $m_{a}=10^{-16}\mathrm{eV}$; Blue, $m_{a}=10^{-18}\mathrm{eV}$.
Fig. 3 indicates that a larger spatial scale of the satellite network is more favorable for detecting axion signals, providing better sensitivity to light axions. Unlike observations using natural pulsars, in our work Eq. (III.4) incorporates the effect of $f$, for light axions with Compton wavelengths $\lambda_{\text{com}}\geq L$, $g_{95\%}$ shows no significant variation on the order of magnitude. A similar behavior can also be seen in Fig. 4, which shows the 95% C.L. upper limits $g_{95\%}$ on the function $g_{a\gamma}\big(L\big)$ for different $m_{a}$. In this figure, the characteristic scale $L$ of the satellite network ranges from $10^{5}\mathrm{km}$ to $10^{9}\mathrm{km}$, covering scales from the Earth–Moon distance to the Jupiter–Sun distance. We find that for lighter axions, the 95% C.L. upper limit is more sensitive to $L$. A larger $L$ leads to a stronger constraint on $g_{a\gamma}$. Likewise, a saturation scale exists: once $L$ exceeds the Compton wavelength $\lambda_{\text{com}}$ of axion, further increasing does not improve the 95% C.L. upper limit.
@@ III.2 Frequentist analysis-GLSP
The GLSP is a spectral analysis technique designed to detect periodic signals in irregularly sampled time series. It employs a least squares fitting procedure to scan the trial frequencies and compute the power spectrum at each frequency. Compared to traditional Fourier transform methods, the GLSP is particularly advantageous for handling unevenly sampled data, making it well suited for astronomical observations [52]. According to [52], the GLSP is expressed as
$P_{LS}\big(f\big)$ $=\frac{1}{\sigma^{2}_{y}\sum_{i}\omega_{i}}\Bigg\{\frac{\big[\sum_{i}\omega_{i}\big(y_{i}-\bar{y}\big)\cos\big(2\pi f\big(t_{i}-\tau\big)\big)\big]^{2}}{\sum_{i}\omega_{i}\cos^{2}\big[2\pi f\big(t_{i}-\tau\big)\big]}$ (III.6)
@@
$\tau=\frac{1}{4\pi f}\arctan\frac{\sum_{i}\omega_{i}\sin\big(4\pi ft_{i}\big)}{\sum_{i}\omega_{i}\cos\big(4\pi ft_{i}\big)},$ (III.7)
@@
$\omega_{i}$ $=\frac{1}{\sigma^{2}_{i}},\quad\bar{y}=\frac{\sum_{i}\omega_{i}y_{i}}{\sum_{i}\omega_{i}},$ (III.8)
@@
A distinctive feature of the window function in ground-based radio telescope observations is that, in order to detect faint pulsar signals, observations are restricted to nighttime. Consequently, the resulting time series is highly non-uniformity sampled. Moreover, due to the annual modulation introduced by the Earth’s orbital motion, the data are not only influenced by the ionospheric interference discussed earlier but also subject to additional deterministic periodic effects [52]. Together, these factors give rise to pronounced sidelobes and spectral aliasing in the Lomb–Scargle periodogram [52], thereby increasing the False Alarm Probability (FAP) and potentially obscuring the true signal peaks. As shown in Fig.4 of [12], the PPTA observed four representative pulsars (J0437–4715, J1643–1224, J1730–2304 and Crab–QUIJOTE) and presented the GLSP of the time series of polarization measurements. Even after applying the IRI ionospheric corrections, the GLSPs of the four pulsars remain highly irregular. The irregular temporal sampling causes the target axion signal to be masked by numerous spurious peaks in the power spectrum. This effect becomes increasingly severe as the observational S/N decreases, and in extreme cases, the signal peak may fall entirely below the significance threshold, rendering it undetectable. The figure also demonstrates that many interference peaks approach the 32% FAP level, some of them even approach the 5% FAP level, substantially complicating the reliable identification of genuine signals.
@@
In contrast, within our proposed framework, once the satellite network is deployed, it can operate in continuous monitoring mode. In the idealized case, the corresponding window function of space-based observations reduces to $W_{{t_{i}}}\big(t\big)=1$, thereby eliminating the non-uniformity inherent in ground-based sampling. To illustrate this, we consider a simplified scenario in which the observed signal contains only a single axion induced frequency component, $m_{a}=10^{-22}\mathrm{eV}$. The transmitting satellites emit periodic pulse signals towards the receiving satellite at a fixed interval $\Delta t=\delta t+\sigma_{t}=\big(1+\xi_{\delta t}\big)\delta t$. In addition, the system is affected by Gaussian white noise, $\phi_{\text{noise}}$. Under these assumptions, the APPA measured at the receiving end can be expressed as
$\Delta\phi_{\text{sim}}=\phi_{a}\mathrm{cos}\big(2\pi\nu+\varphi\big)+\phi_{\text{noise}}.$ (III.9)
@@
Similar to Fig.4 in [12], we consider the case of $\text{S/N}=1$ and a target signal frequency of the axion with mass $m_{a}=10^{-22}\mathrm{eV}$. However, in this scenario, the observation epochs are distributed with high regularity over a continuous 10-year period. The corresponding simulation results are presented in Fig. 5.
@@
Figure 5: Numerical simulation: The GLSP (red) derived from APPA observational data. The simulated signal includes only Gaussian white noise with S/N=1. The vertical dashed line marks the target signal frequency of the axion with mass $m_{a}=10^{-22}\mathrm{eV}$. The horizontal lines indicate FAP levels of 10% (dotdashed), 5% (dashed) and 1% (solid).
@@
If the observational data are of low quality and the target frequency peak is completely masked by interference peaks, the GLSP method becomes ineffective. In such cases, for further information, the reader is referred to [43, 12]. Once the GLSP method successfully extracts the target axion signal frequency $m_{a}$ from the pulsed signals received by the satellite, the APPA sensitivity curve for axion signal detection can be derived using Eq. (B.1) and Eq. (III.9).
We assume that the satellite network possesses a detection sensitivity of $0.01^{\circ}$ for linear polarization angle rotation, and that the instrumental noise is parameterized according to Eq. (III.1) and Eq. (III.2). Under these assumptions, the sensitivity curve can be obtained subject to the constraint of a total observation time of $T_{\text{obs}}\sim\mathcal{O}\big(10^{1}\big)$yr. Note that when the axion mass is sufficiently small ($\mathrm{cos}\big(m_{a}T\big)\simeq 1$, to the left of the dashed line in Fig. 2), Eq. (B.1) reduces to Eq. (B.2), in this limit, the expected value of $\left|\alpha_{o}-\alpha_{s}\right|$ is 0.734; when condition $\mathrm{cos}\big(m_{a}T\big)\simeq 1$ is no longer satisfied (to the right of the dashed line in Fig. 2, we adopt $m_{a}L=0.1$ as the reference criterion), the expectation value of $\big[\big(\alpha_{o}^{2}+\alpha_{s}^{2}\big)-2\alpha_{o}\alpha_{s}\mathrm{cos}\big(m_{a}T\big)\big]^{\frac{1}{2}}$ must be evaluated individually for each value of $m_{a}$. Unlike the likelihood analysis, the spatial scale of the satellite network only determines the critical line at which $\mathrm{cos}\big(m_{a}T\big)\simeq 1$ and does not affect the overall sensitivity curve. Detailed analysis can be found in Appendix B.
@@ IV Conclusion
Inspired by the artificial precision timing array in [2], we investigate the APPA concept, which leverages an artificial satellite network to achieve precise measurements of ultralight scalar axion signals. Compared with PPA data obtained from ground-based observatories, APPA offers several significant advantages: (i) The number of uncertain or uncontrollable physical parameters is greatly reduced, since most parameters can be explicitly determined and flexibly adjusted; (ii) Complex or periodic physical disturbances are largely eliminated, thereby reducing the burden of data analysis and enhancing the quality of the raw observational data; (iii) Due to the ultra precise clocks of satellites and the tunability of pulsed signal parameters, the S/N, periodicity, and long term stability of the measurements are well ensured. Consequently, APPA provides an exceptionally broad detectable frequency range.
To clarify the feasibility and advantages of APPA, we employ two complementary approaches: likelihood analysis and frequentist analysis. For the likelihood analysis, we avoid prohibitively large scale computations by using the convenient expression in Eq. (III.4) from [14, 19], which allows us to efficiently derive the $g_{95\%}$ constraint. Likelihood analysis indicates that larger scales $L$ of satellite networks are more advantageous for detecting low mass axions. However, for axions with $\lambda_{\text{com}}\geq L$, APPA exhibit identical detection limits. Fig. 2 shows that APPA exhibits a detection advantage within $m_{a}\sim\mathcal{O}\big(10^{-22}–10^{-18}\big)$eV, which lies well below the detection upper bound. Compared with traditional ground-based observations, APPA achieves a tighter $g_{95\%}$ bound, highlighting its strong detection potential. For the frequentist approach, we adopt the GLSP, a standard method for detecting periodic sinusoidal signals. MC simulations demonstrate that, even under low S/N conditions, APPA produces GLSP spectra with more pronounced target signal peaks, significantly reduced FAP, and fewer interference peaks compared with conventional observations.
@@
Acknowledgments
We thank helpful conversion with Profs. Szabolcs Marka and Jing Ren. This work is supported by the National Key Research and Development Program of China (Nos. 2023YFC2206200, No.2021YFC2201901), the National Natural Science Foundation of China (No.12375059), and the Project of National Astronomical Observatories, Chinese Academy of Sciences (No.E4TG6601).